OpenStudy (anonymous):

write this trinomial in factored form 4x^2+19x-5

OpenStudy (anonymous):

ah, much better. (4x + 1)(x - 5)

OpenStudy (anonymous):

whoops (4x -1)(x+5)

OpenStudy (anonymous):

x1 = (19+21)/8 x1 = 5 x2 = (19-21)/8 x2 = -1/4 The factored form: (x-x1)(x-x2) = (x-5)(x+ 1/4)